# Enhancing Feature Diversity Boosts Channel-Adaptive Vision Transformers

### v1

@@ Abstract
Multi-Channel Imaging (MCI) contains an array of challenges for encoding useful feature representations not present in traditional images. For example, images from two different satellites may both contain RGB channels, but the remaining channels can be different for each imaging source. Thus, MCI models must support a variety of channel configurations at test time. Recent work has extended traditional visual encoders for MCI, such as Vision Transformers (ViT), by supplementing pixel information with an encoding representing the channel configuration. However, these methods treat each channel equally, i.e., they do not consider the unique properties of each channel type, which can result in needless and potentially harmful redundancies in the learned features. For example, if RGB channels are always present, the other channels can focus on extracting information that cannot be captured by the RGB channels. To this end, we propose DiChaViT, which aims to enhance the diversity in the learned features of MCI-ViT models. This is achieved through a novel channel sampling strategy that encourages the selection of more distinct channel sets for training. Additionally, we employ regularization and initialization techniques to increase the likelihood that new information is learned from each channel. Many of our improvements are architecture agnostic and could be incorporated into new architectures as they are developed. Experiments on both satellite and cell microscopy datasets, CHAMMI, JUMP-CP, and So2Sat, report DiChaViT yields a $1.5-5.0\%$ gain over the state-of-the-art.
@@ 1 Introduction
Most visual encoders assume they are provided with a fixed-channel representation as input (e.g., they take RGB inputs as input at train and test time) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, many applications find a variety of imaging techniques beyond just the traditional RGB channels beneficial. For example, satellite images or sensors onboard a robot often contain an infrared camera in addition to traditional RGB, and microscopes can also host a significant range of potential imaging channels [11, 12, 13, 14, 15, 16, 17]. Thus, Multi-Channel Imaging (MCI) models aim to learn good feature representations from datasets with heterogeneous channels, where the number and type of channels can vary for each input at test time. Training a model that is robust to changes in channel configurations can save time and resources as only a single model needs to be learned, while also helping to prevent overfitting in small datasets through transfer learning [14]. Prior work made MCI methods more robust to missing channels by randomly masking them out during training [18]. As shown in Fig. 1(a) left and (b) top, this results in redundancies being learned across channels during training rather than encoding new information. In effect, by focusing on learning what to do when a channel is missing, the model becomes less able to learn unique and/or more challenging cues within each channel.
@@
To address this limitation, we propose Diverse Channel Vision Transformer (DiChaViT) that aims to better balance the robustness in different MCI configurations, which may cause redundancy, with a need to learning diverse and informative features. First, we include a Channel Diversification Loss (CDL), a regularization term that encourages a special channel token, which represents the presence of a channel in the input data, to be distinct from the other channel tokens. As shown in Fig. 1(a) right, this reduces repeated information across our model’s channels. However, this can still result in similar features being encoded for each image patch. Thus, our Token Diversification Loss (TDL) aims to directly diversify the features learned for each patch token as shown in the bottom of Fig. 1(b) by encouraging that each patch token is orthogonal to the others. Finally, rather than a uniform random channel masking strategy as used in prior work [18, 19], we introduce Diverse Channel Sampling (DCS), in which we select channels based on their dissimilarity, further promoting feature diversity . We observe that promoting a more diverse representation enables each channel to contribute more to the final prediction, leading to a performance boost of up to $5.0$% in downstream MCI tasks. Fig. 2 provides an overview of our approach.
@@
The work that is closest in spirit to ours are methods that are designed to learn disentangled representations [20, 21, 22, 23, 24, 25, 26, 27], e.g., learning features aligned to a given set of attributes [28, 29, 30]. These methods have shown a trade-off between the strength of the disentanglement and the downstream tasks performance [31, 32, 33]. This is due, in part, to the fact that many attributes these methods aim to disentangle are correlated with each other, making it challenging to know what features relate individually to each attribute. However, MCI’s challenge is different in that we focus not only on disentangling features across channels. Instead, we aim to capture redundancies that are required to be robust to missing channels without ignoring the features that may only arise in a subset (or even a single) channel. In other words, in MCI some redundancy is desirable across channels even if we could learn perfectly disentangled representations. In addition, many methods in disentangled representation learning assume the attributes to separate are labeled, but there are no labeled attributes in MCI. Instead, DiChaViT must automatically decide what to capture in multiple channels while still learning important channel-specific information.
@@
Figure 1: Mutual information and attention scores of HPA channels (CHAMMI [14]) after training. (a) The channel tokens in ChannelViT have high mutual information, suggesting there may be some redundancy [34, 35]. In contrast, DiChaViT shows lower mutual information between channels, encouraging each channel token to learn different features. Note we normalize the values and gray out the diagonal for better visualization. (b) We compute attention scores of the [CLS] token to the patch tokens in the penultimate layers and aggregate them by channel. ChannelViT (top) relies on certain channels (e.g., microtubules and nucleus) to make predictions and less on other channels (e.g., protein and er). In contrast, DiChaViT demonstrates more evenly distributed attention scores across channels, suggesting that each channel contributes more to the model’s predictions.
@@ 2 Related Work
Convolutional-based models for multi-channel imaging. Researchers have been developing convolutional-based models to keep pace with the evolving landscape of multi-channel imaging data. Bhattacharyya et al. introduced IRFacExNet, which utilizes depth-wise convolutions to merge channel-wise features from infrared thermal images [36]. Jiang et al. introduced a double-channel CNN that takes into account the correlation between input channels in aerial images [37]. This approach employs a separate sub-network for each group of channels and then performs feature fusion to aggregate features across channels. Siegismund et al. presented DCMIX to work with images with many channels based on imaging blending concepts [38]. While these methods can be used for MCI, they are not designed to work on varying input channels. In a recent study, Chen et al. [14] introduced and adapted channel-adaptive models, such as Depthwise, TemplateMixing [39, 40, 41], and HyperNet [42], which can handle images with varying input channels. These models incorporate adaptive interfaces in their first layers, followed by a shared ConvNeXt backbone [8]. Specifically, the first layer of the ConvNeXt is replaced with a channel-agnostic adapter, while all other layers remain the same. While these methods provide a strong baseline, they find settings where some channels are missing during inference challenging. In our work, we aim to improve MCI model robustness by improving the diversity of learned features.
@@
Vision Transformers for multi-channel imaging. Vision transformers (ViT) [43] have natural advantages when dealing with multiple channels, especially when the number of channels varies. ViTs treat image modeling as sequence-to-sequence problems, allowing them to be flexible in handling different numbers of image tokens. Nguyen et al. introduced variable tokenization and variable aggregation, in which they divided each input channel independently into patches and then aggregated the patch features across channels using learnable queries [44]. Tarasiou et al. proposed TSViT, which incorporates a tokenization scheme and temporal position encodings to process Satellite Image Time Series [45]. In a relevant work, Zhou et al. introduced FAN, a channel reweighting design aimed at adjusting channel features based on the observation that some channels capture more significant information than others [46]. In the medical domain, Hatamizadeh et al. proposed UNETR that utilized a transformer encoder followed by a skip-connected decoder for 3-D medical image segmentation [47]. Recently, Bao et al. proposed ChannelViT that processes each input channel independently via a shared linear projection and incorporates a learnable channel embedding for preserving channel-specific features [18]. In addition, the authors proposed Hierarchical Channel Sampling (HCS), a regularization technique applied to the input channels to boost robustness and reduce training time. ChannelViT outperforms standard ViTs in classification tasks and demonstrates its generalization ability when only a subset of the trained channels is available during inference. In a similar work, Bourriez et al. introduced ChAda-ViT, a channel adaptive attention technique for handling heterogeneous microscope images [15]. However, these methods do not adequately model the unique properties of each channel type, resulting in harmful redundancies, whereas we boost the diversity of features across channels to enhance the robustness of MCI-ViT models.
@@ 3 Encouraging Diverse Representations in multi-channel ViTs
Given a multi-channel image (MCI) $X$ containing channels $c_{i}\in C_{X}$, our goal is to train a model $M$ that takes our input image $X$ as input to make its predictions. Following [15, 18], we consider the MCI setting where $M$ has seen all the channels we expect to see during inference, i.e., $C_{X}\subseteq C_{M}$, and leave the exploration of supporting novel channels during inference to future work. Since we do not know what $C_{X}$ we may see during inference, prior work has focused primarily on exploring methods that are robust to different choices of $C_{X}$ by encouraging $M$ to redundancies across channels (e.g.[14, 15, 18]). Specifically, they begin with a base ViT encoder [48] that uses each channel-specific image patch $p_{i}$ as input. Each image patch is passed through a shared patch projection layer and concatenated with its corresponding channel token $c_{i}$. Hierarchical Channel Sampling (HCS) encourages robustness to missing channels by randomly masking some channels during training to ensure key information can be captured in multiple channels [18]. However, as noted in the Introduction, this can be harmful when $M$ does not balance this repetitive feature learning to also capture distinctive channel-specific information.
As demonstrated in Fig. 1, we address this in DiChaViT by enhancing diversity in our model’s learned representations. First, we use a Channel Diversification Loss (CDL) to learn diverse representations to help prevent feature collapse in the channel tokens (Sec. 3.1). Second, our Token Diversification Loss (TDL) encourages patch tokens to also learn distinct features (Sec. 3.2). Finally, Diverse Channel Sampling (DCS) enhances DiChaViT’s ability to learn new features during training (Sec. 3.3). These components enable our approach to have a better balance between repetitive and channel-specific feature learning (overview in Fig. 2).
@@ 3.1 Enhancing channel token separation
Recall that in Fig. 1(a), learned channel tokens $c_{i}$ show high mutual information, indicating that these tokens are not well-separated. Following [41, 49, 50, 51], we partly mitigate this issue by replacing the random initialization of $c_{i}$ used by prior work [15, 18] with an orthogonal initialization. To further encourage the diversity in the features, we introduce Channel Diversification Loss (CDL) for increased separation between the channel tokens (Fig. 2(a)). The main idea, inspired by ProxyNCA++ loss [52], is to use a learnable vector (i.e., channel anchor) to represent each channel in the input image. During training, we compute the distance between the channel tokens to these anchors. Our objective is to pull channel features toward their corresponding anchors while pushing them away from all other anchors. The anchors prevent channel tokens from collapsing while still allowing for flexibility in learning useful representations.
We start by initializing the channel tokens orthogonally and then applying CDL on these, i.e.,
$\mathcal{L}_{\mathrm{CDL}}=-\log\left(\frac{\exp\left(-d\left(\frac{c_{i}}{\left\|c_{i}\right\|_{2}},\frac{f\left(c_{i}\right)}{\left\|f\left(c_{i}\right)\right\|_{2}}\right)*\frac{1}{t}\right)}{\sum_{f(a)\in A}\exp\left(-d\left(\frac{c_{i}}{\left\|c_{i}\right\|_{2}},\frac{f(a)}{\|f(a)\|_{2}}\right)*\frac{1}{t}\right)}\right),$ (1)
where $f(c_{i})$ is a function that returns a corresponding channel anchor for channel $c_{i}$, and $d(c_{i},f(\cdot))$ is the squared Euclidean distance between channel token $c_{i}$ and an anchor. We denote $A$ as the set of all channel anchor, $t$ as the temperature, and ${\|\cdot\|_{2}}$ as the $L2$-Norm. In Eq. 1, the numerator calculates the distance of a channel token to its anchor, while the denominator computes all these distance pairs of the channel token to all the channel anchors. When the temperature value $t$ is set to 1, we get a standard Softmax function. Lowering the temperature can lead to a more focused and sharp probability distribution, but we found that the results are not very sensitive to the value of $t$. Thus, we simply use a fixed temperature $t$ of $0.07$.
@@ 3.2 Enhancing feature diversity for patch tokens
MCI-ViT models (e.g., ChannelViT, ChAda-ViT) use a shared linear projection to extract features independently from each input channel in the image. This approach has been shown to improve performance and robustness compared to using a separate projection for each channel [18, 15]. With the shared projection, only the common features across channels are retained, while other channel-specific information is filtered out, which helps to reduce overfitting. However, this design could lead also to similar representations being encoded for all the patch tokens. This is not ideal because each patch may contain unique information that could be ignored. To address this issue, we propose Token Diversification Loss (TDL), a regularization applied to the patch token features to enhance the diversity of features learned by each patch in the input image. Fig. 2(b) provides an overview of TDL. Specifically, we enforce an orthogonality constraint on the tokens to ensure that each token is orthogonal to the others. Additionally, we take into account the token type information to differentiate between tokens from the same channels and across channels. The main idea is to make features from different channels more distinct while allowing for a certain level of similarity among features within the same channel.
Let $\mathbf{p}_{i}$ be the input patch at position $i$, and $\mathbf{W_{\mathrm{proj}}}$ be the shared linear projection at the first layer. We denote $\mathbf{t}_{i}=\mathbf{W}_{\mathrm{proj}}\cdot\mathbf{p}_{i}$ be the patch feature token of $\mathbf{p}_{i}$, $T=\bigl{\{}\mathbf{t}_{i}\bigr{\}}_{i=1,2,...}$ is a set containing all patch feature tokens in the input image, and $f(\mathbf{t}_{i})$ is a function that returns the corresponding channel for input patch $\mathbf{p}_{i}$. We devise a unified loss function for each input image as follows.
$\mathcal{L}_{\text{s}}=\frac{1}{N_{s}}\quad\sum_{\begin{subarray}{c}\mathbf{t}_{i},\mathbf{t}_{j}\in T;\text{ }f(\mathbf{t}_{i})=f(\mathbf{t}_{j})\end{subarray}}\left\langle\mathbf{t}_{i},\mathbf{t}_{j}\right\rangle$ (2)
$\mathcal{L}_{\text{d}}=\frac{1}{N_{d}}\quad\sum_{\begin{subarray}{c}\mathbf{t}_{i},\mathbf{t}_{k}\in T;\text{ }f(\mathbf{t}_{i})\neq f(\mathbf{t}_{k})\end{subarray}}\left\langle\mathbf{t}_{i},\mathbf{t}_{k}\right\rangle$ (3)

### v2

@@ Abstract
Multi-Channel Imaging (MCI) contains an array of challenges for encoding useful feature representations not present in traditional images. For example, images from two different satellites may both contain RGB channels, but the remaining channels can be different for each imaging source. Thus, MCI models must support a variety of channel configurations at test time. Recent work has extended traditional visual encoders for MCI, such as Vision Transformers (ViT), by supplementing pixel information with an encoding representing the channel configuration. However, these methods treat each channel equally, i.e., they do not consider the unique properties of each channel type, which can result in needless and potentially harmful redundancies in the learned features. For example, if RGB channels are always present, the other channels can focus on extracting information that cannot be captured by the RGB channels. To this end, we propose DiChaViT, which aims to enhance the diversity in the learned features of MCI-ViT models. This is achieved through a novel channel sampling strategy that encourages the selection of more distinct channel sets for training. Additionally, we employ regularization and initialization techniques to increase the likelihood that new information is learned from each channel. Many of our improvements are architecture agnostic and can be incorporated into new architectures as they are developed. Experiments on both satellite and cell microscopy datasets, CHAMMI, JUMP-CP, and So2Sat, report DiChaViT yields a $1.5-5.0\%$ gain over the state-of-the-art. Our code is publicly available at https://github.com/chaudatascience/diverse_channel_vit.
@@ 1 Introduction
Most visual encoders assume they are provided with a fixed-channel representation as input (e.g., they take RGB inputs as input at train and test time) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, many applications find a variety of imaging techniques beyond just the traditional RGB channels beneficial. For example, satellite images or sensors onboard a robot often contain an infrared camera in addition to traditional RGB, and microscopes can also host a significant range of potential imaging channels [11, 12, 13, 14, 15, 16, 17]. Thus, Multi-Channel Imaging (MCI) models aim to learn good feature representations from datasets with heterogeneous channels, where the number and type of channels can vary for each input at test time. Training a model that is robust to changes in channel configurations can save time and resources as only a single model needs to be learned, while also helping to prevent overfitting in small datasets through transfer learning [14]. Prior work proposed methods to make MCI models robust to missing channels by randomly masking them during training [18]. As shown in Fig. 1(a) left and (b) top, this results in redundancies being learned across channels during training rather than encoding new information. A consequence of this repetition is a model focused on learning strong cues that are easy to identify, making it less capable of learning unique and/or challenging cues within each channel.
@@
To address this limitation, we propose a Diverse Channel Vision Transformer (DiChaViT) that aims to balance the robustness in different MCI configurations, which may cause redundancy, with a need to learning diverse and informative features. First, we include a Channel Diversification Loss (CDL), a regularization term that encourages a special channel token, which represents the presence of a channel in the input data, to be distinct from the other channel tokens. As shown in Fig. 1(a) right, this reduces repeated information across our model’s channels. However, this can still result in similar features being encoded for each image patch. Thus, our Token Diversification Loss (TDL) aims to directly diversify the features learned for each patch token as shown in the bottom of Fig. 1(b) by encouraging that each patch token is orthogonal to the others. Finally, rather than a uniform random channel masking strategy as used in prior work [18, 19], we introduce Diverse Channel Sampling (DCS), in which we select channels based on their dissimilarity, further promoting feature diversity. We observe that promoting a more diverse representation enables each channel to contribute more to the final prediction, leading to a performance boost of up to $5.0$% in downstream MCI tasks. Fig. 2 provides an overview of our approach.
@@
The work that is closest in spirit to ours are methods that are designed to learn disentangled representations [20, 21, 22, 23, 24, 25, 26, 27], e.g., learning features aligned to a given set of attributes [28, 29, 30]. These methods have shown a trade-off between the strength of the disentanglement and the downstream tasks performance [31, 32, 33]. This is due, in part, to the fact that many attributes these methods aim to disentangle are correlated with each other, making it challenging to know what features relate individually to each attribute. However, unlike these tasks, MCI methods do not focus only on disentangling features across channels. Instead, they must capture some redundant information to be robust to missing channels while simultaneously learning features that may only arise in a subset (or even a single) channel. In other words, in MCI some redundancy is desirable across channels even if we could learn perfectly disentangled representations. In addition, many methods in disentangled representation learning assume the attributes to separate are labeled, but there are no labeled attributes in MCI. Instead, DiChaViT must automatically decide what to capture in multiple channels while still learning important channel-specific information.
@@
Figure 1: Comparison of the redundant information learned by different models on the HPA dataset in CHAMMI [14]). (a) Measures the mutual information between the channel tokens, which captured the configuration of channels in an image. Note we gray out the diagonal for better visualization. We find ChannelViT tokens have high mutual information, which suggests significant redundancy exists across channels [34, 35]. In contrast, DiChaViT has little mutual information as each channel is encouraged to learn different features. (b) We compute attention scores of the $\mathrm{[CLS]}$ token to the patch tokens in the penultimate layers and aggregate them by channel. ChannelViT (top) relies on certain channels (e.g., microtubules and nucleus) to make predictions and less on other channels (e.g., protein and er). In contrast, DiChaViT demonstrates more evenly distributed attention scores across channels, suggesting that each channel contributes more to the model’s predictions.
@@ 2 Related Work
Convolutional-based models for multi-channel imaging. Researchers have been developing convolutional-based models to keep pace with the evolving landscape of multi-channel imaging data. Bhattacharyya et al. [36] introduced IRFacExNet, which utilizes depth-wise convolutions to merge channel-wise features from infrared thermal images. Jiang et al. [37] introduced a double-channel CNN that takes into account the correlation between input channels in aerial images. This approach employs a separate sub-network for each group of channels and then performs feature fusion to aggregate features across channels. Siegismund et al. [38] presented DCMIX to work with images with many channels based on imaging blending concepts. While these methods can be used for MCI, they are not designed to work on varying input channels. In a recent study, Chen et al. [14] introduced and adapted channel-adaptive models based Depthwise convolutions, TemplateMixing [39, 40, 41], and HyperNets [42]. These models incorporate their adaptive interface in the first layer of an otherwise shared ConvNeXt model [8]. While these methods provide a strong baseline, they find settings where some channels are missing during inference challenging. In our work, we aim to improve MCI model robustness by improving the diversity of learned features.
@@
Vision Transformers for multi-channel imaging. Vision transformers (ViT) [43] have natural advantages when dealing with multiple channels, especially when the number of channels varies. ViTs treat image modeling as sequence-to-sequence problems, allowing them to be flexible in handling different numbers of image tokens. Nguyen et al. [44] introduced variable tokenization and variable aggregation, in which they divided each input channel independently into patches and then aggregated the patch features across channels using learnable queries. Tarasiou et al. [45] proposed TSViT, which incorporates a tokenization scheme and temporal position encodings to process Satellite Image Time Series. In a relevant work, Zhou et al. [46] introduced FAN, a channel reweighting design aimed at adjusting channel features based on the observation that some channels capture more significant information than others. In the medical domain, Hatamizadeh et al. [47] proposed UNETR that utilized a transformer encoder followed by a skip-connected decoder for 3-D medical image segmentation. Recently, Bao et al. [18] proposed ChannelViT that processes each input channel independently via a shared linear projection and incorporates a learnable channel embedding for preserving channel-specific features. In addition, the authors proposed Hierarchical Channel Sampling (HCS), a regularization technique applied to the input channels to boost robustness and reduce training time. ChannelViT outperforms standard ViTs in classification tasks and demonstrates its generalization ability when only a subset of the trained channels is available during inference. In a similar work, Bourriez et al. [15] introduced ChAda-ViT, a channel adaptive attention technique for handling heterogeneous microscope images. However, these methods do not adequately model the unique properties of each channel type, resulting in harmful redundancies, whereas we boost the diversity of features across channels to enhance the robustness of MCI-ViT models.
@@ 3 Encouraging Diverse Representations in multi-channel ViTs
Given a multi-channel image (MCI) $X$ containing channels $c_{i}\in C_{X}$, our goal is to train a model $M$ that takes our input image $X$ as input to make its predictions. Following [15, 18], we consider the MCI setting where $M$ has seen all the channels we expect to see during inference, i.e., $C_{X}\subseteq C_{M}$. We leave the exploration of handling novel channels during inference for future work, as it presents significant challenges, including establishing meaningful connections between existing and new channels, and identifying informative channel weights in the presence of domain shifts. In our setting, since we do not know what $C_{X}$ we may see during inference, prior work has focused primarily on exploring methods that are robust to different choices of $C_{X}$ by encouraging $M$ to redundancies across channels (e.g.[14, 15, 18]). Specifically, they begin with a base ViT encoder [48] that uses each channel-specific image patch $\mathrm{p}_{i}$ as input. Each image patch is passed through a shared patch projection layer and concatenated with its corresponding channel token $\mathrm{ch}_{i}$. Hierarchical Channel Sampling (HCS) [18] encourages robustness to missing channels by randomly masking some channels during training to ensure key information can be captured in multiple channels. However, as noted in the Introduction, this can be harmful when $M$ does not balance this repetitive feature learning to also capture distinctive channel-specific information.
As illustrated in Fig. 1, DiChaViT aims to better balance repetitive and distinct feature learning through three major components. First, we use a Channel Diversification Loss (CDL) to learn diverse representations to help prevent feature collapse in the channel tokens (Sec. 3.1). Second, our Token Diversification Loss (TDL) encourages patch tokens to also learn distinct features (Sec. 3.2). Finally, Diverse Channel Sampling (DCS) promotes robustness to missing channels while also encouraging that new features are also learned during training (Sec. 3.3). These components enable our approach to balance repetitive and channel-specific feature learning (overview in Fig. 2).
@@ 3.1 Enhancing channel token separation
Recall that in Fig. 1(a), learned channel tokens $\mathrm{ch}_{i}$ from prior work show high mutual information, indicating these tokens are not well-separated. Following [41, 49, 50, 51], we partly mitigate this issue by replacing the random initialization of $\mathrm{ch}_{i}$ used by prior work [15, 18] with an orthogonal initialization. To further encourage the diversity in the features, we introduce Channel Diversification Loss (CDL) for increased separation between the channel tokens (Fig. 2(a)). Inspired by ProxyNCA++ [52], the idea is to use a learnable vector (i.e., an orthogonally initialized channel anchor) to represent each channel in the input image during training. We promote diversity in the channel tokens by pulling channel features toward their corresponding anchors while pushing them away from all other anchors. A key benefit of this approach is that the anchors prevent channel tokens from collapsing while still allowing for flexibility in learning useful representations.
Formally, we denote $A$ as the set of all channel anchors, $t_{\mathrm{CDL}}$ as the temperature, and ${\|\cdot\|_{2}}$ as the $L2$-Norm. We start by initializing the channel tokens $\mathrm{ch}_{i}$ and their channel anchors orthogonally. Then, we apply CDL as follows:
$\mathcal{L}_{\mathrm{CDL}}=-\log\left(\frac{\exp\left(-d\left(\frac{\mathrm{ch}_{i}}{\left\|\mathrm{ch}_{i}\right\|_{2}},\frac{g\left(\mathrm{ch}_{i}\right)}{\left\|g\left(\mathrm{ch}_{i}\right)\right\|_{2}}\right)\cdot\frac{1}{t_{\mathrm{CDL}}}\right)}{\sum_{g(a)\in A}\exp\left(-d\left(\frac{\mathrm{ch}_{i}}{\left\|\mathrm{ch}_{i}\right\|_{2}},\frac{g(a)}{\|g(a)\|_{2}}\right)\cdot\frac{1}{t_{\mathrm{CDL}}}\right)}\right),$ (1)
where $g(\mathrm{ch}_{i})$ is a function that returns a corresponding channel anchor for channel token $\mathrm{ch}_{i}$, and $d(\mathrm{ch}_{i},g(\cdot))$ is the squared Euclidean distance between channel token $\mathrm{ch}_{i}$ and an anchor. In Eq. 1, the numerator calculates the distance of a channel token to its anchor, while the denominator computes all these distance pairs of the channel token to all the channel anchors. When the temperature value $t_{\mathrm{CDL}}$ is set to $1$, we get a standard $\mathrm{Softmax}$ function. Lowering the temperature can lead to a more focused and sharp probability distribution, but we found that the results are not very sensitive to the value of $t_{\mathrm{CDL}}$. Thus, we simply use a fixed temperature $t_{\mathrm{CDL}}$ of $1/14\approx 0.07$.
@@ 3.2 Enhancing feature diversity for patch tokens
MCI-ViT models like ChannelViT [18], ChAda-ViT [15] use a shared linear projection to extract features independently from each input channel in the image rather than using separate projections for each channel. With the shared projection, only the common features across channels are retained, while other channel-specific information is filtered out, which helps to reduce overfitting. However, this design can also produce similar representations for all patch tokens. This is not ideal because each patch may contain unique information that would be ignored. In our approach, we also leverage this shared projection, but we enhance it with Token Diversification Loss (TDL), a regularization applied to the patch token features to enhance the diversity of features learned by each patch in the input image (see Fig. 2(b) for an overview). Specifically, we enforce an orthogonality constraint on the tokens to ensure that each token is orthogonal to the others. Additionally, we take into account the token type information to differentiate between tokens from the same channels and across channels. The main idea is to make features from different channels more distinct while allowing for a certain level of similarity among features within the same channel.
Let $\mathbf{p}_{i}$ be the input patch at position $i$, and $\mathbf{W_{\mathrm{proj}}}$ be the shared linear projection at the first layer. We denote $\mathbf{t}_{i}=\mathbf{W}_{\mathrm{proj}}\cdot\mathbf{p}_{i}$ as the patch feature token of $\mathbf{p}_{i}$, $T=\bigl{\{}\mathbf{t}_{i}\bigr{\}}_{i=1,2,...}$ as the set containing all patch feature tokens in the input image, and $h(\mathbf{t}_{i})$ as a function that returns the corresponding channel for input patch $\mathbf{p}_{i}$. We devise a unified loss function for each input image as follows:
$\mathcal{L}_{\text{s}}=\frac{1}{N_{s}}\quad\sum_{\begin{subarray}{c}\mathbf{t}_{i},\mathbf{t}_{j}\in T;\text{ }h(\mathbf{t}_{i})=h(\mathbf{t}_{j})\end{subarray}}\left\langle\mathbf{t}_{i},\mathbf{t}_{j}\right\rangle$ (2)
$\mathcal{L}_{\text{d}}=\frac{1}{N_{d}}\quad\sum_{\begin{subarray}{c}\mathbf{t}_{i},\mathbf{t}_{k}\in T;\text{ }h(\mathbf{t}_{i})\neq h(\mathbf{t}_{k})\end{subarray}}\left\langle\mathbf{t}_{i},\mathbf{t}_{k}\right\rangle$ (3)